# Learning Dynamic Bayesian Networks from Data: Foundations, First Principles and Numerical Comparisons

### v1

@@ 5.3 Enforcing Acyclicity
The primary challenge in solving optimization problems on DAGs stems from the exponential size of the acyclicity constraint. A well-known method to ensure acyclicity involves using cycle elimination constraints, which were originally introduced in the context of the Traveling Salesman Problem (TSP) in [11]. Supposing that the set of all cycles is denoted by $\mathcal{C}$, these constraints often take the form
@@
where $e_{i,j}$ denote binary decision variables that indicate which edges are present in the directed graph. These constraints may be complemented by different score functions to complete the optimization problem leading to dag recovery. This can then lead to different types of problems, some of which are linear [51, 34, 45], some quadratic [56]. Furthermore, this method of cycle elimination is also typically augmented with a cutting plane method [48, 56].
Another method for acyclicity enforcement is derived from a well-known combinatorial optimization problem called linear ordering (LO) [28]. In the LO problem, we aim to find ”the best” permutations, which may be further constrained. In the case a directed acyclic graphs, these permutations correspond to the placements of edges and since the basis has only quadratic cardinality, the number of constraints is limited. The cycles are then excluded by imposing LO constraints. A perceived drawback of this approach is the neccessity for a quadratic cost function [57, 28].
The third method for eliminating cycles involves enforcing constraints to ensure the nodes adhere to a topological order. A topological order is a linear arrangement of the nodes in a graph such that an arc $\left(j,k\right)$ exists only if node $j$ precedes node $k$ in this order. The discrete decision variables, indexed by the node and placement in the topological order determine the graph. It has been reported that in some cases this approach can lead to polynomial time learning [60].
Recently, an alternative approach based on layered networks has been proposed [57]. The concept of layering forbids the placement of arcs between layers in a given direction. The problem of finding a layered graph is defined by the number of layers and the minimal number of layers for a given DAG is unique. This contrasts the topological order method described in the previous paragraph, which can have a possitive influence on the construction of the branch-and-bound tree [57].
@@
The algorithm NOTEARS [72] and DYNOTEARS[50] uses the following functional constraint in a continuous optimization algorithm to enforce the DAG structure of the graph,
@@
In [70] they introduce a different constraint term that also enforces the DAG constraint, but appears to have better numerical stability, for small $\mu>0$:
@@
In the procedure NO BEARS [41] the spectral radius is used to define the presence of a DAG constraint on the adjacency graph. Certain numerical approximations make this relatively feasible, despite the high complexity and nondifferentiability of the spectral radius of a matrix.
Finally, [71] present DAGS with NOCURL, which obviates the need for an explicit functional constraint by solving:
@@
6 Methods for Learning Structure and Parameters in DBNs
@@ 6.1.1 Constraint Based
Under the assumptions of causal sufficiency (no hidden confounders) and faithfulness, classical algorithms developed by Spirtes et al. [62] have been proven to estimate the DAG without exhaustive enumeration of possible structures (which is impossible in interesting cases). The Peter-Clark (PC) algorithm is a method to retrieve the skeleton and directions of the edges, relying on an empirical hypothesis test of Conditional Independence (CI) for each pair of variables given a subset of other variables. It starts from a complete undirected graph and deletes sequentially edges based on these CI relations. PCMCI [55] is adapted to time-series datasets and works for lagged links (causes precede effects). It operates in two stages: 1/ PC testing which identifies a potential set of parents with high probabilities for each variable $X_{j}^{t}$. 2/ using these parents as conditions for the momentary conditional independence (MCI) to address the false positives and test all variable pairs. Statistical tests ParCorr, GPDC, and CMI are used in both steps. PCMCI+ extends PCMCI to include contemporaneous links [54].
@@ 6.1.2 Score Based:
The Integer Programming based [3] uses the local score (BDeu, BGe, DiscreteLL, DiscreteBIC, DiscreteAIC, GaussianLL, GaussianBIC, GaussianAIC, GaussianL0) to optimize the network, amortizing its evaluation, thus obviating the need to compute parameters to compute the score. This algorithm was later relased as GOBNILP [14] (Globally Optimal Bayesian Network learning using Integer Linear Programming). GOBNILP finds the network with the highest BDeu score under the constraint that the underlying structure can be represented as a DAG. For every node in the graph $v$ and every possible parent set $W$, binary variable $I(W\rightarrow v)$ is created. The optimization criterion is then sum over all possible vertices and all possible parent sets, where the BDeu score for the selected parent set of every node is considered, i.e.,
@@
The constraints are then of two types. First, each vertex needs to have only a single parent set, which for node $v$ formulates as
@@
The second constraint requires that there are no cycles in the graph. This is imposed by cluster constraints, which require that there must be $1$ node with no parents for any set of nodes. As there are exponentially many such sets of nodes, the optimization problem is solved, and if a cycle is in the final solution, the cluster constraint that prohibits the found cycle is added. Such computation is iterated until a DAG is found, which also ensured that the optimal model is found.
@@
GFlowNet (GFN) for structure learning [17] consists of approximating the posterior instead of finding a single DAG, to reduce uncertainty over models. They construct the sample DAG from the posterior as a sequential decision problem by starting from an empty graph and adding one edge at a time. The GFN environment is similar to Reinforcement Learning where the states are different graphs, each associated with a reward which is the score of that structure. They define a terminal state $s_{f}$ to which every connected state is called complete. The actions taken are edge adding (no edge reversal or removal). In addition, they define a mask that prevents having cycles in the graphs. GFN’s goal is to model the whole distribution proportional to the rewards. It also borrows from Markov chain literature, using forward and backward transition probabilities, $P_{\theta}(s^{\prime}/s)$ and $P_{B}(s/s^{\prime})$ in the loss function that satisfies the detailed balance condition:
@@
where $R(s)$ is the reward function of state $s$. Extending GFN to DBN required changing the scoring function BDe and BGe adequately and changing the mask used before to also take into account the stationarity assumption (transitions are invariant in time) and to be a block upper triangular matrix (no edges going from time slice $t+1$ to $t$).
This has been recently extended in [18] for sampling the structure and weights simultaneously using recent developments of expanding the GFN environment to continuous variables [39].
@@
Monte Carlo methods are classical for solving difficult statistical problems, and have been a popular choice for learning the structure and parameters of a DBN. There are two prominent Monte Carlo methods in the literature that developed the foundations and have been seminal in the development of structure learning algorithms. These include the work (1128 citations as of this writing)  [24] as well as (2344) [63], who developed the popular MMHC, a max-min hill climb (MMHC) procedure.
In the numerical experiments, we use MMHC from the package bnstruct [23].
@@
We use [38], a more recent development. It uses order based structure sampling and at the same restricts the search space using conditional independence tests. Performance of the method is generally the strongest performer for difficult problems.
@@ 6.1.3 One Shot Linear SEMs:
There are two prominent procedures that represent one shot learning of LSEMs. The two are based on integer and continuous based optimization. LSEMs indeed uniquely presents the opportunity for continuous optimization methods, and as such presents the possibility of scaling the estimation procedure, at the cost of theoretical guarantees of global convergence.
@@
The mixed integer-linear program defined in [45] is presented here:
@@
We begin by presenting the general algorithm introduced in [50] which followed the well cited [72]. In this paper, they consider the transition dynamics of $X(t)$ can be expressed using the SEM:
@@
6.2 Novel Modifications of Existing Methods
In developing the work for this paper, a few natural developments of existing algorithms, that wouldn’t be worthwhile to appear independently, arose. We present each of these methods and describe them
One Shot Structure-Parameter Consistent Frequentist
We propose a modified variant of (46). In this case, we introduce positive and negative weights, and require a lower bound for the weights. Thus, if the edge is active, the weights are forced to be bounded away from zero. This is based on two motivations:
  1. 1.
    In principle, a structure being correctly identified should correspond to the weights associated with any active edge to be nonzero. Thus a search for the structure fitting the data well should be expected to have weights that would reject a null hypothesis of zero.
  2. 2.
    In the literature on sparsity ($\|\cdot\|_{0}$) constrained optimization, e.g. [4], it can be seen that a necessary (but not sufficient) condition for optimality is an $L$-stationarity condition that implies that, effectively, the indices $\mathcal{I}(\theta^{*})=\mathop{supp}(\theta^{*})$ are such that $\theta_{i},i\in\mathcal{I}(\theta^{*})$ must be bounded away from zero a distance corresponding the Lipschitz constant of the gradient and the gradient vector components corresponding to the components of $\theta^{*}$ that are zero.
$\begin{array}[]{rl}\min\limits_{\left(\begin{array}[]{c}E_{W^{+}},E_{W^{-}},E_{A^{+}},E_{A^{-}}\\ W^{+},W^{-},A^{+},A^{-}\end{array}\right)}&\mathbf{E}((E_{W^{+}},E_{W^{-}},E_{A^{+}},E_{A^{-}},W^{+},W^{-},A^{+},A^{-}))\\ &+\lambda_{W^{+}}\|E_{W^{+}}\|_{0}+\lambda_{W^{-}}\|E_{W^{-}}\|_{0}+\lambda_{A^{+}}\|E_{A^{+}}\|_{0}+\lambda_{A^{-}}\|E_{A^{-}}\|_{0}\\ &:=\sum\limits_{m=1}^{M}\sum\limits_{t=1}^{T}\sum\limits_{i=1}^{d}\left([X_{m,t}]_{i}-\sum\limits_{j=1}^{d}[E_{W^{+}}]_{l,j,i}W^{+}_{j,i}[X_{m,t}]_{j}-\sum\limits_{j=1}^{d}[E_{W^{-}}]_{l,j,i}W^{-}_{j,i}[X_{m,t}]_{j}\right.\\ &\left.-\sum\limits_{l=1}^{\min\{p,t\}}\sum\limits_{j=1}^{n}[E_{A^{+}}]_{l,j,i}A^{+}_{l,j,i}[X_{m,t-l}]_{j}-\sum\limits_{l=1}^{\min\{p,t\}}\sum\limits_{j=1}^{n}[E_{A^{-}}]_{l,j,i}A^{-}_{l,j,i}[X_{m,t-l}]_{j}\right)^{2}\\ &+\lambda_{W^{+}}\sum\limits_{i,j}[E_{W^{+}}]_{i,j}+\lambda_{W^{-}}\sum\limits_{i,j}[E_{W^{-}}]_{i,j}\\ &+\lambda_{A^{+}}\sum\limits_{l,i,j}[E_{A^{+}}]_{l,i,j}+\lambda_{A^{-}}\sum\limits_{l,i,j}[E_{A^{-}}]_{l,i,j}\\ \text{s.t. }&W^{+}\geq b_{W},\,W^{-}\leq-b_{W}\\ &A^{+}\geq b_{A},\,A^{-}\leq-b_{A}\\ &E_{W^{+}}+E_{W^{-}}\leq 1,\,E_{A^{+}}+E_{A^{-}}\leq 1\\ &DAG\left(E_{W^{+}}+E_{W^{-}}\right),\\ &E_{W^{+}},E_{W^{-}}\in\left[\{0,1\}^{d^{2}}\right]\\ &E_{A^{+}},E_{A^{-}}\in\left[\{0,1\}^{d^{2}}\right]^{p}\\ &W^{+},W^{-}\in\mathbb{R}^{d\times d},\,A^{+},A^{-}\in\mathbb{R}^{p\times d\times d}\end{array}$ (46)
where $b_{W},b_{A}>0$ are lower bounds on the magnitude of these weights.
@@ 7 Numerical Results
Note: this is a work on progress, and the numerical results reported here are exploratory
The synthetic datasets were generated following causalLens [40]. We set the maximum lag to 1 and the graph complexity to 30, corresponding to complex causal graphs. The algorithms used are Dynotears with hyperparameters $lambda\_w$ and $lambda\_a$ equal to 0.05, and a small $w\_threshold$ of 0.01. For structure identification in tables 4 and 5, we only compare the structure of the algorithms, so the binary adjacency matrix (rather than the weighted one) is taken from Dynotears. For GOBNILP, the algorithm only supports IID data. To use it, we run it twice on the data in the first time slice to get the prior network, then on the two first slices to get the transition network. For the MCMC, we use iterative MCMC followed by order MCMC from BiDAG package, to sample the MAP DAG. We set $alpha$ to , $alphainit$ to 0.01 and change $hardlimit$, limit on the size of parent sets , according to the number of variables per experiment. For PCMCI+, we choose an $pc\_alpha$ of 0.01 and use ParCorr as the conditional independence test (which assumes univariate, continuous variables with linear dependencies and Gaussian noise). We further correct the p-values by False Discovery Rate control with an $alpha\_level$ of 0.01. And finally we use Max Min Hill Climbing algorithm from bnstruct package using the BIC score. The parameters of the one shot frequentist ILP approach (see (46)) were $b_{W}=b_{A}=0.1$. The regularization parameters were $\lambda_{A^{+}}=\lambda_{A^{-}}=\lambda_{W^{+}}=\lambda_{W^{-}}=0.05$.
@@
We can consider three main purposes for which DBNs may be used for, and so we perform tests comparing the learners for these three criteria in an appropriate manner. In addition, we report on the time of execution, and present results across the scale of small and medium covariate dimension problems.
@@
    Ground Truth Graph Identification One of the primary goals of using BN and DBN models for fitting various time-varying phenomena is causal discovery and causal inference. This amounts to being able to accurately reconstruct the graph from a noisy realization of the ground truth. Indeed under the causal identifiability assumption given above, the relative success by which a learner is able to compute this ground truth graph is, understandably so, a central for evaluating DBN learners in the literature.
@@
: Favorable Regime for Identification: This corresponds to $NT\gg n$, in which case, the more generally well-developed methods are able to identify the ground truth graph. We shall take:
$(n,N,T)\in\{(3,30,10),(5,50,50),(10,100,200)\}$ (47)
@@
$(n,N,T)\in\{(3,5,10),(5,10,20),(10,20,40),(20,40,50),(30,60,100)\}$ (48)

### v2

@@ 4.3 Enforcing Acyclicity
The primary challenge in solving optimization problems on DAGs stems from the exponential size of the acyclicity constraint. A well-known method to ensure acyclicity involves using cycle elimination constraints, which were originally introduced in the context of the Traveling Salesman Problem (TSP) in [chvatal10]. Supposing that the set of all cycles is denoted by $\mathcal{C}$, these constraints often take the form
@@
where $e_{i,j}$ denote binary decision variables that indicate which edges are present in the directed graph. These constraints may be complemented by different score functions to complete the optimization problem leading to dag recovery. This can then lead to different types of problems, some of which are linear [park17, jaakkola2010learning, manzour2021integer], some quadratic [rytir2024exdag]. Furthermore, this method of cycle elimination is also typically augmented with a cutting plane method [Nemhauser91, rytir2024exdag].
Another method for acyclicity enforcement is derived from a well-known combinatorial optimization problem called linear ordering (LO) [gro85]. In the LO problem, we aim to find ”the best” permutations, which may be further constrained. In the case a directed acyclic graphs, these permutations correspond to the placements of edges and since the basis has only quadratic cardinality, the number of constraints is limited. The cycles are then excluded by imposing LO constraints. A perceived drawback of this approach is the neccessity for a quadratic cost function [seoh2020solving, gro85].
The third method for eliminating cycles involves enforcing constraints to ensure the nodes adhere to a topological order. A topological order is a linear arrangement of the nodes in a graph such that an arc $\left(j,k\right)$ exists only if node $j$ precedes node $k$ in this order. The discrete decision variables, indexed by the node and placement in the topological order determine the graph. It has been reported that in some cases this approach can lead to polynomial time learning [Shojaie10].
Recently, an alternative approach based on layered networks has been proposed [seoh2020solving]. The concept of layering forbids the placement of arcs between layers in a given direction. The problem of finding a layered graph is defined by the number of layers and the minimal number of layers for a given DAG is unique. This contrasts the topological order method described in the previous paragraph, which can have a possitive influence on the construction of the branch-and-bound tree [seoh2020solving].
@@
The algorithm NOTEARS [zheng2018dags] and DYNOTEARS[pamfil2020dynotears] uses the following functional constraint in a continuous optimization algorithm to enforce the DAG structure of the graph,
@@
In [yu2019dag] they introduce a different constraint term that also enforces the DAG constraint, but appears to have better numerical stability, for small $\mu>0$:
@@
In the procedure NO BEARS [lee2019scaling] the spectral radius is used to define the presence of a DAG constraint on the adjacency graph. Certain numerical approximations make this relatively feasible, despite the high complexity and nondifferentiability of the spectral radius of a matrix.
Finally, [yu2021dags] present DAGS with NOCURL, which obviates the need for an explicit functional constraint by solving:
@@
5 Structure and Parameter Learning Algorithms
5.1 General Principles and Considerations
Here we present a number of important points of consideration that can be observed from looking at the literature at successful attempts at representation as far as inference and learning. With the distinctions described below, we are able to properly identify and characterize existing structure-weight learners, as well as suggest extensions to fill in the natural empty places in any observed taxonomy.
Structure Learning
Parameter Learning
[…]
Frequentist and Bayesian
We shall use the frequentist versus Bayesian distinction to indicate a point estimate based on the optimization of a loss function or criterion, and a probabilistic model, implemented with sampling, that obtains a posterior distribution of the structure and weights given the data, respectively. A frequentist estimate is given as a complete specific structure encoding and a specific value for the parameters. It is generally expected, or at least sought, that the relationships that the graph identifies between various nodes representing the random variables is statistically significant. This presumption often becomes unrealistic in practice, and obtaining an appropriately scaled statistically significant entire network, that is, with all significant edges, is typically unavailable in real settings. However, since DBNs are generative rather than discriminative, this is often not a practical concern, as they are a component in an overall statistical modeling pipeline.
An effective and commonly used technique is to employ mixtures of a finite set of structures, see [friedman2000being]. This provides flexibility and the transparent uncertainty in the model, without having to mix through the entire combinatorial space defining possible structure.
For instance, consider a particular point estimate of a structure and set of parameters. However, consider that the set of parameters is close to zero, and moreover, that is so close so as to include zero in a, e.g., 95% confidence interval. In this case, it is clear that this implies that the presence of this edge itself in the graph is suspect, that is, not implied by the data to statistical significance.
Hierarchical and One-Shot Methods
In general one can consider most learning methods to be hierarchical in the sense of first learning the structure, and with an amortized structure estimating or sampling the weights. The use of SEMs defined by adjacency matrices including both structure and weights simultaneously introduced what can be referred to as a one-shot approach (we remark the interestingly similar recent popularity of one-shot methods for neural architecture search that simultaneously perform parameter learning [guo2020single]).
In this case, a point estimate is obtained for both the structure and the weights simultaneously by solving an appropriate optimization problem that fits both of these as decision variables to the data. To this end there are two approaches we see in the literature. In [manzour2021integer] an IP (for BNs, readily adapted to DBNs) is presented that treats the structure as binary variables encoding the activation of edge links in the graph and the parameters as separate variables, and solves the challenging nonlinear mixed IP (relaxation into conic programs was considered in [kucukyavuz2023consistent]). Alternatively, the recent work DYNOTEARS [atanackovic2024dyngfn] presented a gradient based method for solving the structure-parameter learning as a purely continuous optimization problem for weight matrices in the graph. Enforcing sparsity is done to encourage accurate ground truth graph identification.
We make one additional remark going back to hierarchical approaches. Note that one can consider that the frequentist-Bayesian distinction can be applied to present a taxonomy of methods. As a curious example, many Bayesian scoring methods, e.g. the IP method [BARTLETT2017258], can be considered hybrid frequentist-Bayesian. This is because implicitly the grading is done with a Bayesian parameter model, but a point estimate, that is one unique structure, is returned. Methodologically, we see that the advantages of a hierarchical is an offline calculation of scoring that permits the use of simple and powerful off the shelf commercial grade IP solvers, and the disadvantage is the conceptual contradiction of applying a frequentist mindset to learning structure with Bayesian models as weights. However, one can easily mitigate this in practice by sampling from multiple structures, as weighted in frequency by their respective marginal likelihoods. Regardless, theoretically, in the asymptotic regime, consistency can still be maintained with all approaches and variations thereof, however [koller2009probabilistic].
5.2 Methods for Learning Structure and Parameters in DBNs
@@ 5.2.1 Constraint Based
Under the assumptions of causal sufficiency (no hidden confounders) and faithfulness, classical algorithms developed by Spirtes et al. [causaldiscov_PC_FCI] have been proven to estimate the DAG without exhaustive enumeration of possible structures (which is impossible in interesting cases). The Peter-Clark (PC) algorithm is a method to retrieve the skeleton and directions of the edges, relying on an empirical hypothesis test of Conditional Independence (CI) for each pair of variables given a subset of other variables. It starts from a complete undirected graph and deletes sequentially edges based on these CI relations. PCMCI [pcmci] is adapted to time-series datasets and works for lagged links (causes precede effects). It operates two stages: 1) PC testing which identifies a potential set of parents with high probabilities for each variable $X_{j}^{t}$. 2) using these parents as conditions for the momentary conditional independence (MCI) to address the false positives and test all variable pairs. Statistical tests ParCorr, GPDC, and CMI are used in both steps. PCMCI+ extends PCMCI to include contemporaneous links [runge2022discovering].
@@ 5.2.2 Score Based:
The Integer Programming based [BARTLETT2017258] uses the local score (BDeu, BGe, DiscreteLL, DiscreteBIC, DiscreteAIC, GaussianLL, GaussianBIC, GaussianAIC, GaussianL0) to optimize the network, amortizing its evaluation, thus obviating the need to compute parameters to compute the score. This algorithm was later released as GOBNILP [cussens2020gobnilp] (Globally Optimal Bayesian Network learning using Integer Linear Programming). GOBNILP finds the network with the highest score under the constraint that the underlying structure can be represented as a DAG. For every node in the graph $v$ and every possible parent set $W$, binary variable $I(W\rightarrow v)$ is created. The optimization criterion is then sum over all possible vertices and all possible parent sets, where the BDeu score (we use BDeu score here as an example) for the selected parent set of every node is considered, i.e.,
@@
The constraints are then of two types. First, each vertex needs to have a single parent set, which for node $v$ formulates as
@@
The second constraint requires that there are no cycles in the graph. This is imposed by cluster constraints, which require that there must be $1$ node with no parents for any set of nodes. As there are exponentially many such sets of nodes, the optimization problem is solved, and if a cycle is in the final solution, the cluster constraint that prohibits the found cycle is added. Such computation is iterated until a DAG is found, which ensures that the optimal model is obtained.
@@
GFlowNet (GFN) for structure learning [pmlr-v180-deleu22a] consists of approximating the posterior instead of finding a single DAG, to reduce uncertainty over models. They construct the sample DAG from the posterior as a sequential decision problem by starting from an empty graph and adding one edge at a time. The GFN environment is similar to previous approaches in Reinforcement Learning, where the states are different graphs, each associated with a reward, which is in this case the score of that structure. They define a terminal state $s_{f}$ to which every connected state is called complete. The possible actions are limited to edge adding (no edge reversal nor removal). In addition, they define a mask that prevents having cycles in the graphs. GFN’s goal is to model the whole distribution proportional to the rewards. It also borrows from Markov chain literature, using forward and backward transition probabilities, $P_{\theta}(s^{\prime}/s)$ and $P_{B}(s/s^{\prime})$, in the loss function that satisfies the detailed balance condition:
@@
where $R(s)$ is the reward function of state $s$. Extending GFN to DBN required changing the scoring functions BDe and BGe adequately and changing the static GFN’s mask to take into account the stationarity assumption (transitions are invariant in time) and present it as a block upper triangular matrix, to ensure that no edges are going from time slice $t+1$ to $t$.
This has been recently extended in [deleu2024joint] for sampling the structure and weights simultaneously using recent developments of expanding the GFN environment to continuous variables [lahlou2023theory].
@@
Monte Carlo methods are classical for solving difficult statistical problems, and have been a popular choice for learning the structure and parameters of a DBN. There are two prominent Monte Carlo methods in the literature that developed the foundations and have been seminal in the development of structure learning algorithms. These include the work (1128 citations as of this writing)  [friedman2000being] as well as (2344) [tsamardinos2006max], who developed the popular MMHC, a max-min hill climb (MMHC) procedure.
In the numerical experiments, we use DMMHC from the package dbnR, which uses a variation of dynamic max-min hill climbing [trabelsi2013new], proper for leaning DBNs.
@@
We use [kuipers2022efficient], a more recent development. It uses order based structure sampling and at the same time restricts the search space using conditional independence tests. This method, implemented in the BiDAG package, can be viewed in a way as a hybrid based method, combining both constraint and score based methodologies. Performance of the method is generally the strongest for smaller datasets.
@@ 5.2.3 One Shot Linear SEMs:
There are two prominent procedures that represent one shot learning of LSEMs. The two are based on integer and continuous based optimization. LSEMs indeed uniquely present the opportunity for continuous optimization methods, and as such present the possibility of scaling the estimation procedure, at the cost of theoretical guarantees of global convergence.
@@
The mixed integer-linear program defined in [manzour2021integer] is presented here:
@@
We begin by presenting the general algorithm introduced in [pamfil2020dynotears] which followed the well cited [zheng2018dags]. In this paper, they consider the transition dynamics of $X(t)$ can be expressed using the SEM:
@@ 6 Numerical Results
The synthetic datasets were generated following causaLens [lawrence2021data] and the DBN data simulation code from the DBNclass package (available at https://github.com/cbg-ethz/DBNclass) from [psuter]. For causaLens, we set the maximum lag to 1 and the graph complexity to 30, representing complex causal graphs.
  • GFN: the prior network is learned using the original GFN algorithm from [pmlr-v180-deleu22a], and we learn the transition network using the modified version of the code with the BGe score for the DBN. The hyperparameters were kept as default which can be found on the GitHub repository https://github.com/tristandeleu/jax-dag-gflownet and we only change the number of iterations according to each experiment.
  • Dynotears: for small datasets we do a grid search for the hyperparameters over the following values: $lambda\_a$ and $lambda\_w\in\{0.01,0.05,0.1,0.2,0.5\}$ and $w\_threshold\in\{0.0,0.01,0.02,0.05\}$. For larger datasets ($(20,400,400)$, $(30,60,100)$, $(30,600,500)$) that require more time, we only use the algorithm with hyperparameters $lambda\_w$ and $lambda\_a$ equal to $0.1$, and a small $w\_threshold$ of $0.01$.
  • GOBNILP: the algorithm only supports IID data. To use it, we run it twice on the data: on the first time slice to get the prior network, then on the two first slices to get the transition network.
  • MCMC: we use iterative MCMC followed by order MCMC from BiDAG package, to sample the MAP DAG. We start by a grid search for the small datasets where we choose the values $alpha\in\{0.01,0.05,0.1,0.2,0.3,0.4,0.5\}$ and we set the $hardlimit$ on the size of parent sets as $2n$ where $n$ is the number of covariates. For the large datasets, we restrict $alpha$ to $0.01$ and change $hardlimit$ according to the number of variables per experiment ($hardlimit=20$ for $n=20$ and $hardlimit=10$ for $n=30$).
  • PCMCI+: we choose a range of $pc\_alpha\in\{0.01,0.05,0.1,0.2,0.5,0.7\}$ for small datasets, and starting from $n=20$ and $n=30$, we fix $pc\_alpha$ at 0.01. We choose ParCorr as the conditional independence test (which assumes univariate, continuous variables with linear dependencies and Gaussian noise). We further correct the p-values by False Discovery Rate control with an $alpha\_level$ of $0.01$.
  • HC-dbnR: we keep default parameters (see ttps://github.com/dkesada/dbnR).
% The parameters of the one shot frequentist ILP approach (see (LABEL:eq:iposmod)) were \$b˙W = b˙A = 0.1\$. The regularization parameters were \$λ˙A^+ = λ˙A^- = λ˙W^+ = λ˙W^- = 0.05\$.
@@
We can consider two main purposes for which DBNs may be used, and so we perform tests comparing the learners for both criteria in an appropriate manner. In addition, we present results across the scale of small and medium covariate dimension problems. % In addition, we report on the time of execution, and present results across the scale of small and medium covariate dimension problems.
@@
    Ground Truth Graph Identification One of the primary goals of using BN and DBN models for fitting various time-varying phenomena is causal discovery and causal inference. This amounts to being able to accurately reconstruct the graph from a noisy realization of the ground truth. Indeed under the causal identifiability assumption given above, the relative success by which a learner is able to compute this ground truth graph is, understandably so, a central for evaluating DBN learners in the literature. %
  3. 3.
    Prediction and Forecasting Unique to DBNs, given that it is a time series based model, we are able to consider its accuracy as far as forecasting the future. This can be done by synthetically creating a ground truth DBN, then add one more time slice with the values of the variables estimated using the ground truth. This will be compared to the forward evolution, i.e., a forecast from the learned model.
@@
: Favorable Regime for Identification: This corresponds to \$NT≫n\$, in which case, the more generally well-developed methods are able to identify the ground truth graph. We shall take:
$(n,N,T)\in\{(3,30,10),(5,50,50),(10,100,200),(20,400,400),(30,600,500)\}$ (47)
@@
[…]
For each experiment in these data regimes, we generate 10 datasets and run the algorithms. We report the mean and standard deviation across our datasets in the following sections. For the algorithms where we performed the grid search, we keep the best results based on both SHD and AUROC equally. The chosen parameters for each experiment are listed in Table LABEL:table:params_dynotears and LABEL:table:params_pcmci. As for MCMC, the best performing parameter was \$alpha = 0.01\$ for all the experiments.